OpenStudy (ambermarie151):

The function c(f) =5/9 (f – 32) allows you to convert degrees Fahrenheit to degrees Celsius. Find the inverse of the function so that you can convert degrees Celsius back to degrees Fahrenheit. A. f(c) =5/9 c – 32 B. f(c) =9/5 c + 32 C. f(c) = 9/5c – 32 D. f(c) = 5/9c + 32

OpenStudy (ambermarie151):

@trojanpoem

OpenStudy (danjs):

change the farhrenheigt and celsius around, and solve for celsius// f = 5/9 * (c - 32) (9/5)*f + 32 = c so c^(-1)(f) = (9/5)*f + 32